OpenStudy (anonymous):

If an object is thrown striaght up at an initial velocity of 9.8 m/s, how long will it take to fall back to the ground?

OpenStudy (anonymous):

Is that all you were given? No time or distance, I had a problem similiar to this, but I am not sure there is a way to solve with only the velocity.

OpenStudy (anonymous):

Look at only the first half of the projectile motion. Upward motion: Vi = +9.8 m/s Vf = 0 m/s a = -9.8 m/s^2 Vf = Vi + at --> (Vf-Vi)/a = t Since projectile motion is symmetric, the answer is 2*t.